An Army camp consists of 800 soldiers and their canteen has food provision that last for 60 days. If 400 more soldiers arrive at the camp after 12 days, for how many days will the food last?
a)52
b)48
c)28
d)32

Answer by ankor@dixie-net.com(22740)

An Army camp consists of 800 soldiers and their canteen has food provision that last for 60 days.
If 400 more soldiers arrive at the camp after 12 days, for how many days will the food last?
:
there are 800 * 60 = 48000 man-days of food available
After 12 days the original 800 men have consumed 12*800 = 9600 man-days of food
therefore 48000 - 9600 = 38400 man-days remain, but there are 1200 soldiers now
let d = no. of days the food will last now
1200d = 38400
d = 38400/1200
d = 32 days